OpenStudy (anonymous):

Find the slope ,if it exists of the line containing the pair of points. (-9,-7) and (-19,-16).

OpenStudy (anonymous):

(y2-y1)/(x2-x1)

OpenStudy (anonymous):

what is that the answer

OpenStudy (amistre64):

that is the formula you can use to FIND the answer ...

OpenStudy (anonymous):

how would i get theanswer

OpenStudy (amistre64):

i have a tendency to put the wrong parts in the wrong places tho, so i go a different route

OpenStudy (amistre64):

subtract one point from the other to start out with

OpenStudy (amistre64):

(-19,-16) - ( -9, -7) ------------ x, y ; then stack y/x for the slope

OpenStudy (anonymous):

10 and 9

OpenStudy (amistre64):

good; x=10, y=9 ; so... y/x = 9/10 = slope

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

the "formula" goes thru the same process, but you have to make sure your placing the values in correctly; and i never do :)

OpenStudy (anonymous):

well thanks anyhow